Dinitrogen pentoxide is added to an evacuated rigid container at 25 degrees Celsius. Given that the initial pressure is 1.00 atm and at equilibrium the pressure is 0.500 atm, determine the final pressure in the reaction vessel at equilibrium.
Answer #1

the reaction carried out is

2 N2O5 <---> 4 NO2 + O2

using the ICE table

initial pressure of N2O5 , NO2 , O2 are 1 , 0 , 0

change in pressure of N2O5 , NO2 , O2 are -2x , +4x , +x

equilibrium pressure of N2O5 , NO2 , O2 are 1-2x , 4x , x

we can see that

equilibrium pressure of N2O5 = 1-2x = 0.5

1-2x = 0.5

2x = 0.5

x = 0.25

now

final pressure in the reaction vessel = pN205 + pNO2 + pO2

final pressure in the reaction vessel = 1-2x + 4x + x

final pressure in the reaction vessel = 1 + 3x

final pressure in the reaction vessel = 1.75 atm